OpenStudy (anonymous):

is y-2 and -2y the same thing?

jimthompson5910 (jim_thompson5910):

if you meant to say y(-2) and -2y, then they are the same y(-2) means y times -2 -2y means -2 times y you can multiply two numbers in any order which is why y(-2) = -2y

jimthompson5910 (jim_thompson5910):

however, y-2 is not the same as -2y

OpenStudy (anonymous):

i have to find the inverse of the function f(x) = 5x+2

jimthompson5910 (jim_thompson5910):

f(x) = 5x+2 y = 5x+2 ... replace f(x) with y x = 5y + 2 ... swap x and y now solve for y

OpenStudy (anonymous):

this is what i did but apparently is wrong... f(x) = 5x+2 y = 5x + 2 y-2 = 5x x = (y-2) / 5 x = (2-y) / 5 y = (2-x)/5 f^-1(x) = 2-x/5

OpenStudy (ineedhelpfast12):

no they are not

OpenStudy (anonymous):

they are not wrong?

OpenStudy (anonymous):

nope they're different

jimthompson5910 (jim_thompson5910):

y-2 is not the same as y-2

OpenStudy (anonymous):

then how exactly do i do this? i'm not sure how to find the inverse of that equation

jimthompson5910 (jim_thompson5910):

when you got to x = (y-2) / 5, you did it correctly once you got to x = (2-y) / 5, you made your first mistake

jimthompson5910 (jim_thompson5910):

I would stop at x = (y-2) / 5 and call the inverse \[\Large f^{-1}(x) = \frac{x-2}{5}\]

OpenStudy (anonymous):

oh i see what i did wrong!! thank you for the help!